OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x f(x)=x^2-3 g(x)=sq3+x

OpenStudy (anonymous):

Let's do the first one f(g(x)...do you know how to do composition of functions?

OpenStudy (anonymous):

can we find the g(f(x)) first

OpenStudy (anonymous):

that the one i don't understand

OpenStudy (anonymous):

Is that\[\sqrt{3+x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So in that it is g(f(x)...we take the g(x) above and where ever we see an x we substitute the value for f(x)...so what would we end up with?

OpenStudy (anonymous):

would the answer be g(f(x))=sq root 3+(x^2-3)=sq x^2=x

OpenStudy (anonymous):

You seem to understand it quiet well..good job. So how about the other equation.

OpenStudy (anonymous):

would the other one be f(g(x))=(sq root 3+x)^2-3=(3+x)-3=x

OpenStudy (anonymous):

Winner winner chicken dinner..great job!!!!

OpenStudy (anonymous):

thank you i just wanted to make sure i did it right